OpenStudy (angelwings996):

Simplify the complex fraction (1/x - 1/y^2) / (1/x^2y - 1/xy^2)

OpenStudy (angelwings996):

\[\frac{ \frac{ 1 }{ x } - \frac{ 1 }{ y ^{2} }}{ \frac{ 1 }{ x ^{2}y } - \frac{ 1 }{ xy ^{2} } }\]

OpenStudy (kropot72):

Look at the denominator. I/x is a factor of both terms in the denominator. So factorise the denominator and you are nearly there!

OpenStudy (angelwings996):

So just factor them for right now ?

OpenStudy (kropot72):

Yes. Can you do it?

OpenStudy (angelwings996):

Yes hold on a second, Im doing the math right now

OpenStudy (kropot72):

Sure :)

OpenStudy (angelwings996):

@kropot72 I'm not really getting anything :/

OpenStudy (kropot72):

Putting the numerator terms over a common denominator: \[\frac{1}{x}-\frac{1}{y ^{2}}=\frac{y ^{2}-x}{xy ^{2}}.............(1)\] Putting the denominator terms over a common denominator: \[\frac{1}{x ^{2}y}-\frac{1}{xy ^{2}}=\frac{y-x}{x ^{2}y ^{2}}...........(2)\] Dividing (1) by (2) gives: \[\frac{xy ^{2}-x ^{2}}{y-x}\]

OpenStudy (angelwings996):

Is that the answer or is there more simplifying to do?

OpenStudy (kropot72):

I can't simplify it further.

OpenStudy (angelwings996):

Well in the back of the book it says the answer is x so maybe they simplified the y-x ?

OpenStudy (kropot72):

Are you sure the question has been copied correctly?

OpenStudy (angelwings996):

Im pretty sure it is.... hold on, ill see if another textbook has the same question or something

OpenStudy (kropot72):

Thank you. I really want to help!

OpenStudy (angelwings996):

Okay so I looked on the online book and it says it is 1/xy not 1/x

OpenStudy (angelwings996):

And thanks, I appreciate the help (: @kropot72

OpenStudy (kropot72):

You're welcome. Sorry, I have reached a dead end :(

OpenStudy (angelwings996):

It's fine, I think I can figure it out now, thank you for your help!

OpenStudy (kropot72):

You're welcome :)